THEORETICAL STUDIES OF TIME-RESOLVED PHOTOELECTRON SPECTRA OF IBr^{−}

\underline{SAMANTHA HORVATH}, ANNE B. McCOY, and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

In the present study, we examine the time-resolved photoelectron spectra of IBr^{−}. In the photodetachment studies performed by Sanov and co-workers and Lineberger and co-workers,^{a} the anionic species, prepared in its electronic ground state (^{2}Σ_{1/2}^{+}), is excited to either its Ã′ (^{2}Π_{3/2}) or B̃ (^{2}Σ_{1/2}^{+}) excited state, before electron photodetachment and dissociation on the C̃ (^{1}Π_{1}) or higher-lying excited states of the neutral species, respectively. In this work, we use the electronic structure program Columbus to calculate the six lowest electronic states of IBr^{−} and the ten lowest Σ states of IBr at the MR-SO-CISD/aug-cc-pVDZ level of theory/basis, using relativistic core potentials for I and Br. Experimentally determined electronic states of IBr are also used.^{b} Vibrational eigenstates for these electronic states are calculated in a discrete variable representation,^{c} and propagation of the thermally populated X̃-state vibrational wave functions on either the Ã′ or B̃ electronic states of the anion is performed using a Lanczos scheme. We then take time-dependent overlaps between these propagated states and the vibrational eigenstates of the neutral surface. Results for IBr^{−} show good agreement with the experimental time-resolved spectra. Extensions to IBr^{−}(CO_{2})_{n} (n < 2) will also be discussed.

aR. Mabbs, K. Pichugin, and A. Sanov, J. Chem. Phys., 2005, 122, 174305; Leonid Sheps, Elisa M. Miller, and W. C. Lineberger (private communication).

bE. Wrede, S. Laubach, S. Schulenburg, A. Brown, E. R. Wouters, A. J. Orr-Ewing, and M. N. R. Ashfold, J. Chem. Phys., 2001, 114(6), 2629.

cD. T. Colbert and W. H. Miller, J. Chem. Phys., 1992, 96(3), 1982.